### v1

# SuperCarver: Texture-Consistent 3D Geometry Super-Resolution for High-Fidelity Surface Detail Generation

Abstract
Traditional production workflow of high-precision 3D mesh assets necessitates a cumbersome and laborious process of manual sculpting by specialized modelers. The recent years have witnessed remarkable advances in AI-empowered 3D content creation. However, although the latest state-of-the-arts are already capable of generating plausible structures and intricate appearances from images or text prompts, the actual mesh surfaces are typically over-smoothing and lack geometric details. This paper introduces SuperCarver, a 3D geometry super-resolution framework particularly tailored for adding texture-consistent surface details to given coarse meshes. Technically, we start by rendering the original textured mesh into the image domain from multiple viewpoints. To achieve geometric detail generation, we develop a deterministic prior-guided normal diffusion model fine-tuned on a carefully curated dataset of paired low-poly and high-poly normal renderings. To optimize mesh structures from potentially imperfect normal map predictions, we design a simple yet effective noise-resistant inverse rendering scheme based on distance field deformation. Extensive experiments show that SuperCarver generates realistic and expressive surface details as depicted by specific texture appearances, making it a powerful tool for automatically upgrading massive outdated low-quality assets and shortening the iteration cycle of high-quality mesh production in practical applications.
Figure 1: Given a 3D textured model lacking geometric details as input, SuperCarver generates realistic and expressive 3D surface details consistent with its texture appearance within 1 $\sim$ 2 minutes, significantly enhancing the production efficiency of high-quality mesh assets. (Row 1: The input 3D textured meshes. Rows 2&4: The original coarse surfaces and normal maps. Rows 3&5: The resulting high-fidelity mesh surfaces and normal maps after our geometry super-resolution.)
footnotetext: Corresponding author.
1 Introduction
3D modeling plays a key role in industrial design and digital entertainment fields, such as movies, video games, robotics, and virtual/augmented reality. In practice, the accuracy and quality of 3D mesh models directly influence specific downstream performances. However, creating high-quality mesh assets depends heavily on the manual efforts of specialized 3D modelers, which can be cumbersome and laborious.
Generally, the complete industrial 3D modeling pipeline can be broadly divided into three sequential stages in terms of the geometric complexity of the corresponding mesh surface: 1) Given initial concept designs, e.g., 2D sketch drawings, a mid-poly mesh is established to roughly replicate the overall spatial structure. 2) Subsequently, surface sculpting is performed to produce highly expressive surface details on a high-poly mesh, which typically consists of million-scale faces. 3) To ensure the efficiency of the rendering engine, a low-poly mesh with significantly simplified faces is created based on the high-poly mesh, followed by UV unwrapping, detail (such as normal/bump map) baking, and texture painting. Among the above procedures, the effects of high-poly geometry sculpting greatly determine the ultimate graphics quality and visual realism, but it accounts for a large portion of the entire manual workload. Therefore, there is an urgent need to particularly investigate the automated generation of high-fidelity surface details.
The recent advancements of deep learning-based generative modeling have led to remarkable strides in 3D content generation [51]. Restricted by insufficient 3D training data, early zero-shot architectures such as DreamFields [37] and DreamFusion [78] resort to distilling prior knowledge from pre-trained vision-language multi-modal models [84] or 2D image diffusion models [87]. However, such per-shape optimization process can typically suffer from inefficiency and inter-view inconsistency problems. Since the emergence of large-scale mesh datasets [23, 22], feed-forward 3D generation frameworks such as LRM [35] and InstantMesh [117] have received increasing attention, which can achieve end-to-end training and finish inference within several seconds. To further enhance generation quality without significantly compromising efficiency, another family of research such as SyncDreamer [58], Wonder3D [59], and Unique3D [113] proposes to apply multi-view diffusion models to synthesize 2D novel views and then reconstruct the target 3D mesh via iterative optimization procedures. In addition, analogous to 2D image diffusion, native 3D latent diffusion architectures such as CLAY [127] also demonstrate tremendous potential. Overall, the current state-of-the-art 3D generation methods are already capable of producing plausible geometric structures and intricate texture appearances. However, the actual mesh surfaces tend to be over-smoothing and lack geometric details, as shown in Figure 2. Moreover, it is worth noting that existing 3D generation approaches are designed for creating new 3D assets from fundamental creative elements, i.e., reasoning the underlying 3D structures from 2D images or text prompts. This means that they cannot be directly applied on a given 3D model for further enhancement, because there is no guarantee that the original structural characteristics could still be well preserved. In addition to image/text-to-3D generation paradigms, there also exist several lines of relevant studies, such as point cloud upsampling [123, 81], reference-based [33, 10] or text-guided [85, 29] 3D editing, geometric (e.g., depth/normal) map [2, 41, 27] or PBR material [104, 107] estimation. However, there is still a lack of specialized explorations for enhancing the surface details of specifically given 3D models.
In this paper, we present SuperCarver, a novel geometry super-resolution framework operating on a given 3D model for generating realistic and expressive surface details as depicted by its texture appearance. Considering the difficulties of directly learning 3D structural details, we choose normal maps, a 2.5D representation format, to efficiently bridge 2D and 3D domains. Accordingly, we start by rendering the input textured mesh into image space, which facilitates leveraging the power and pre-trained knowledge of image diffusion architectures. Still, accurately and robustly predicting normal maps through diffusion faces several challenges. On the one hand, normals are a form of sensitive and vulnerable geometric property, i.e., moderate numerical variations can lead to obvious damage to the underlying surface. Besides, without sufficient priors, normals predicted in image space often fail to align with the actual surface structures of given meshes. On the other hand, due to the inherent stochasticity of diffusion models, it is inevitable that normal predictions vary across different inferences. To address these problems, inspired by previous attempts to stabilize [45, 116] and accelerate [124, 108] image diffusion mechanisms, we adopt a chain of deterministic interpolations from rendered coarse normal maps to desired high-precision normal maps. In the meantime, necessary conditioning signals are incorporated by extracting semantic and geometric priors from rendered color images and depth maps.
After enhancing normal maps from multiple viewpoints, it is straightforward to perform gradient-based mesh updating [43]. Since optimizing from potentially imperfect multi-view predicted normal maps is prone to causing severe surface degradation, we customize a simple yet effective noise-resistant inverse rendering approach based on distance field deformation to deduce the corresponding surface details.
We conduct comprehensive performance evaluations and comparisons on a manually selected subset of textured mesh models from Objaverse [23] and ABO [21] datasets, which strongly demonstrate the effectiveness of our approach. We also verify that SuperCarver can be applied on off-the-shelf image/text-to-3D frameworks to further enhance the geometry quality of their generation results.
Figure 2: The generation results of the current state-of-the-arts, (a) InstantMesh [117] & (b) CLAY [127], typically lack high-fidelity surface details. The corresponding geometry enhancement results of our proposed SuperCarver are shown in (c) & (d).
2 Related Work
2.1 3D Content Generation
3D content generation [93, 51, 54] has always been a highly valuable yet rather challenging task. Unfortunately, the limited data volume and diversity of classic 3D shape repositories [115, 8] largely restrict the advancement of this field. In order to bypass the reliance on large-scale 3D training data, zero-shot schemes [37, 69, 71] are designed to constrain the text-prompted 3D generation process using the off-the-shelf language-image prior of CLIP [84]. With the advent of pre-trained 2D image diffusion models [34, 95, 87], score distillation [78, 105, 68, 53] is investigated to inject text guidance into the optimization procedure of various neural field representations [70, 42]. Subsequent researches further make continuous progress in improving generation quality and/or speed [11, 110, 62, 19, 100, 120, 49, 83], enabling 3D editing [85, 97], and supplementing image conditioning capabilities [66, 98, 80, 92]. Nevertheless, the distillation-based generation paradigm necessitates sophisticated processing to deal with inefficiency and inter-view inconsistency.
Thanks to the significant increase in the scale of 3D data [23, 22], it becomes promising to directly train feed-forward 3D generation networks. The pioneering work of LRM [35] builds an end-to-end regression framework to predict a triplane [7] NeRF representation, while Instant3D [46] develops a two-stage framework that begins with generating the “view-compiled” image grid from texts. Numerous follow-up studies continue to improve performances from different perspectives [103], such as shifting to the 3DGS [42] representation (LGM [99]), replacing transformers with convolutional architectures (CRM [111]), and adding more geometric supervisions (InstantMesh [117]).
To further boost the generation quality while maintaining satisfactory efficiency, viewpoint-conditioned 2D diffusion [56, 55] is explored to synthesize novel 2D views, followed by iterative optimization for sparse-view 3D reconstruction. However, lifting imperfect multi-view 2D predictions to 3D can be non-trivial and usually problematic. To ensure multi-view consistency, SyncDreamer [58] designs synchronized multi-view noise prediction by constructing an intermediate 3D feature volume. In Zero123++ [91], multiple images are tiled into a single frame to model the joint distribution of all pre-defined views. Wonder3D [59] builds a new multi-view cross-domain 2D diffusion pipeline controlled by a domain switcher to generate multi-view normal maps and color images, after which explicit 3D geometry is extracted via normal fusion. Era3D [47] incorporates explicit regression of camera parameters and designs more efficient and compact multi-view attention. Unique3D [113] constructs a powerful image-to-3D framework with comprehensive improvements in multi-view generation and upscaling of color images and normal maps, together with an efficient multi-stage mesh reconstruction algorithm.
Analogous to 2D diffusion for image generation, another promising direction is to develop native 3D diffusion architectures [44, 114, 125, 50, 127] for conditional 3D generation with different design choices of latent space representations. Additionally, auto-regressive generation of triangle sequences [94, 12, 112, 101, 15] and UV-domain diffusion [118, 26] driven by regular geometry image representations [30, 128, 129] also demonstrate new potential.
2.2 Inverse Rendering and Mesh Optimization
Inverse rendering aims at reasoning the underlying physical attributes of a target 3D scene from its 2D image observations. Different 3D scene parameters such as geometry, texture, material, and lighting can be optimized via backpropagating the gradients w.r.t. the 2D rendering outputs [72, 32]. Earlier research investigates various differentiable renderers [60, 40, 57, 13]. Later studies [5, 6, 96, 130, 126, 38] devote to learning neural field representations. For gradient-based mesh optimization, the most straightforward way is to explicitly perform surface deformation from pre-defined mesh templates [106, 31]. Nevertheless, directly optimizing the mesh vertex positions is prone to surface defect and degeneracy, thus requiring careful initialization, regularization, and remeshing [73, 76]. Another choice is to employ occupancy and distance fields [67, 13, 77] for implicit surface modeling. Since traditional iso-surface extraction algorithms [61, 25, 39, 74] cannot propagate gradients, various differentiable schemes [16, 18] are designed to operate on discrete occupancy grids (DMC [52]), continuous signed distance fields (MeshSDF [86]), or deformable tetrahedrons (DefTet [28], DMTet [89]). Moreover, NJF [1] depicts the mappings between pairs of source and target triangles via Jacobian matrices. FlexiCubes [90] incorporates carefully-chosen learnable parameters into [74] for achieving much more flexible local adjustment of mesh vertices.
2.3 3D Geometry Enhancement
Instead of directly generating textured meshes from images or text prompts, another line of research aims to enhance the geometric surface of given 3D models.
Similar to image super-resolution [109], point cloud upsampling [123, 121, 81, 48, 79, 82, 65, 129] can produce a denser set of spatial points uniformly located on the underlying 3D surface from sparse and possibly noisy inputs. However, these approaches are rarely capable of generating new geometric details. When a high-quality source mesh is given as the reference, detail transfer [4, 33] can synthesize similar geometric patterns on the target surface. More recent works [17, 9, 10] further apply such an examplar-based paradigm on coarse voxels for few-shot 3D shape detailization. Text-guided mesh editing frameworks [29, 3, 19] facilitate flexible and convenient geometry manipulation, but face challenges in accurate and fine-grained control. Additionally, PBR material generation [14, 122, 104, 107] typically involves estimating normal/bump maps, which can be potentially exploited for refinement purposes.
3 Proposed Method
Figure 3: The overall workflow of our proposed SuperCarver for texture-consistent geometry super-resolution. Given an input 3D textured model, we start by rendering color, depth, and normal maps, which are fed into a customized deterministic prior-guided diffusion architecture to generate detail-boosted normal maps. To perform gradient-based mesh refinement from potentially imperfect multi-view normal map predictions, we design a noise-resistant distance field deformation process to drive the optimization of high-fidelity 3D surface details.
3.1 Overall Workflow
Given a 3D textured model $\mathcal{M}$ with vertices $\mathcal{V}$ and faces $\mathcal{F}$, we aim at generating high-fidelity surface details to produce the updated geometric mesh structure $\mathcal{M}^{u}=(\mathcal{V}^{u},\mathcal{F}^{u})$. As illustrated in Figure 3, we start by rendering $\mathcal{M}$ from certain specified viewpoint to obtain the corresponding color image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, normal map $\mathcal{N}\in\mathbb{R}^{H\times W\times 3}$, and depth map $\mathcal{D}\in\mathbb{R}^{H\times W}$. We achieve geometry super-resolution with a two-stage pipeline. In the first stage, we construct a deterministic prior-guided 2D diffusion architecture to predict a high-precision normal map $\mathcal{H}\in\mathbb{R}^{H\times W\times 3}$. By performing inferences from multiple different viewpoints, we can acquire a set of multi-view normal map predictions $\{\mathcal{H}_{k}\}_{k=1}^{K}$. In the second stage, we design a noise-resistant inverse normal rendering approach to iteratively optimize the original mesh surface with geometric constraints, where we perform distance field deformation by learning grid offsets. In what follows, we will detail the two stages.
3.2 Deterministic Prior-Guided Normal Diffusion
We resort to the most popular text-to-image latent diffusion architecture [87] to leverage its strong generalization power and adapt it to generate texture-consistent detail-rich normal maps. Some previous studies [58, 91, 59, 113] resort to joint multi-view diffusion to mitigate inconsistency across views. Still, such a modeling paradigm also restricts the flexibility in adjusting the number and positions of viewpoints if users have personalized needs to designate specific surface areas for geometric detail carving, which are common in practical scenarios. Moreover, since in our setting there is no need to infer invisible novel views from limited visual observations, inter-view inconsistency will not be a major problem. Thus, we choose the more flexible single-view working mode.
To achieve this goal, the most straightforward solution is to denoise the noisy latent representation of the target normal map $\mathcal{H}$ while incorporating the color image $\mathcal{I}$ into the denoising U-Net. Thus, the forward process is given by:
~$\widetilde{\mathcal{H}}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{H}}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\quad t=\{1,...,T\},$ (1)
where the noise schedule term $\bar{\alpha}_{t}$ controls the transformation to the standard Gaussian distribution $\epsilon$, $T$ is the number of time steps, $\widetilde{\mathcal{H}}$ denotes the latent representation of $\mathcal{H}$ after VAE’s encoder, and ~$\widetilde{\mathcal{H}}_{t}$ is the noisy feature map.
Given the denoising network $\mu_{\theta}$, the optimization objective can be formulated as:
$\mathcal{L}_{\theta}=\mathbb{E}_{\epsilon,\widetilde{\mathcal{I}},\mathcal{H},\mathbf{c},t}\big{[}\|\epsilon-\mu_{\theta}(\widetilde{\mathcal{I}},\widetilde{\mathcal{H}}_{t},\mathbf{c},t)\|^{2}_{2}\big{]},$ (2)
where $\mathbf{c}$ is a vectorized embedding originally extracted from texts. The visual cues can be injected via concatenating ~$\widetilde{\mathcal{H}}_{t}$ with the encoded latent feature map (denoted as $\widetilde{\mathcal{I}}$) of $\mathcal{I}$.
Deterministic Diffusion. In fact, the above implementation together with various existing approaches [59, 27, 113] can be regarded as cross-domain (RGB2Normal) image translation, while our special setting motivates us to directly refine the rendered coarse normal map $\mathcal{N}$ to generate the desired high-precision normal map $\mathcal{H}$. To this end, we investigate a deterministic diffusion mechanism, whose forward process can be formulated as:
~$\widetilde{\mathcal{H}}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{H}}+\sqrt{1-\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{N}},\quad t=\{1,...,T\},$ (3)
where we directly replace the Gaussian noise $\epsilon$ (in Eq. (1)) with the latent representation $\widetilde{\mathcal{N}}$ of the rendered normal map $\mathcal{N}$ after VAE’s encoder.
Empirically, eliminating the inherent stochastic nature in this way naturally ensures deterministic outputs and avoids the fluctuations of diffused normal maps across views, thus implicitly enhancing inter-view consistency and stabilizing the subsequent multi-view inverse rendering process.
Prior Extraction. To boost the accuracy and robustness of normal diffusion, we further extract semantic and geometric priors as conditioning signals.
For the rendered color image $\mathcal{I}$, we extract its DINO [75] feature $\widetilde{\mathcal{Q}}$ and also obtain its latent representation $\widetilde{\mathcal{I}}$ through VAE’s encoder. The spatial resolution of $\widetilde{\mathcal{Q}}$ is pre-upscaled by several auxiliary convolutional layers. Besides, we apply CLIP’s image encoder to extract the vectorized embedding $\tilde{\mathbf{c}}$ for cross-attention within U-Net’s transformer blocks. For the rendered depth map $\mathcal{D}$, we treat its latent representation $\widetilde{\mathcal{D}}$ as the geometric prior. Then we concatenate $\widetilde{\mathcal{Q}}$, $\widetilde{\mathcal{I}}$, and $\widetilde{\mathcal{D}}$ to deduce the resulting conditioning signal $\widetilde{\mathcal{S}}$.
Thus, we can achieve prior-guided diffusion by concatenating $\widetilde{\mathcal{S}}$ with the noisy latent representation ~$\widetilde{\mathcal{H}}_{t}$. The corresponding optimization objective can be formulated as:
$\mathcal{L}_{\theta}=\mathbb{E}_{\widetilde{\mathcal{N}},\widetilde{\mathcal{S}},\mathcal{H},\tilde{\mathbf{c}},t}\big{[}\|\widetilde{\mathcal{N}}-\mu_{\theta}(\widetilde{\mathcal{S}},\widetilde{\mathcal{H}}_{t},\tilde{\mathbf{c}},t)\|^{2}_{2}\big{]}.$ (4)
Empirically, we introduce another parameterization form of v-prediction [88] to train the denoising U-Net, whose input channels and cross-attention dimensions are adapted to match with our prior-conditioned latent representations and CLIP-encoded image embeddings.
3.3 Noise-Resistant Inverse Normal Rendering
Figure 4: Illustration of applying our proposed SuperCarver for further 3D surface detail generation on textured meshes generated from (a) InstantMesh [117], (b) Unique3D [113], and (c) CLAY [127].
When finishing the training of our normal diffusion model, we can render multi-view images from different viewpoints $\{\mathbf{v}_{k}\}_{k=1}^{K}$, and perform inference to correspondingly predict a set of detail-rich normal maps $\{\mathcal{H}_{k}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{K}$.
To transfer the underlying 3D structural patterns implied by the predicted 2D normal maps to the input mesh surface, the common practice is to perform gradient-based mesh refinement through inverse rendering. However, it is observed that updating the mesh geometry from potentially imperfect normal map predictions is prone to causing obvious surface degradation, even with adaptations of various existing state-of-the-art (explicit [76] or implicit [90]) mesh optimization approaches. Hence, we are strongly motivated to develop a highly noise-resistant scheme better suited to our setting of adding new details to a given coarse surface.
We introduce a simple yet effective distance field deformation approach, working in a “value-fixing, grid-moving” manner. Specifically, assuming that the original mesh $\mathcal{M}=(\mathcal{V},\mathcal{F})$ has already been normalized into the unit sphere, we start by constructing uniform 3D grid points $\mathcal{G}\in\mathbb{R}^{r\times r\times r\times 3}$ within $[-1,1]^{3}$ and computing per-grid signed distance values $\mathcal{X}\in\mathbb{R}^{r\times r\times r}$. Different from common ways of implementing deformable distance field representations via only updating $\mathcal{X}$ or jointly updating $\mathcal{G}$ and $\mathcal{X}$, we choose a subtle strategy that fixes distance values $\mathcal{X}$ while separately learning zero-initialized offsets $\mathcal{O}\in\mathbb{R}^{r\times r\times r\times 3}$ to move initial grid positions $\mathcal{G}$. In comparison, standard distance field refinement requires higher resolutions of fixed uniform grids; otherwise, detailed structures cannot be captured even with accurate distance values. Additionally, fixing pre-computed distance values hinders the optimization process from drastically changing the original mesh structure even with obviously erroneous normal prediction in certain areas.
To achieve this, we employ a differentiable implementation of a classic iso-surface extraction algorithm DMC [74], which we denote as $\Psi(\cdot;\cdot)$, for producing an updated mesh $\mathcal{M}^{u}=(\mathcal{V}^{u},\mathcal{F}^{u})$. Then, a differentiable rasterizer $\mathcal{R}(\cdot;\cdot)$ is integrated to render the correspondingly updated normal maps $\{\mathcal{H}_{k}^{u}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{K}$ w.r.t. every viewpoint $\mathbf{v}_{k}$. The whole processing procedure can be formulated as:
$\mathcal{G}^{u}=\mathcal{G}+\tau\cdot\sigma(\mathcal{O}),$ (5)
$\mathcal{H}_{k}^{u}=\mathcal{R}(\Psi(\mathcal{G}^{u};\mathcal{X});\mathbf{v}_{k}),$ (6)
where we bound the range of grid offset values by applying the hyperbolic tangent (Tanh) function $\sigma(\cdot)$ multiplied by a positive scalar $\tau$, then the deformed grid points $\mathcal{G}^{u}$ and fixed distance values $\mathcal{X}$ are used for the subsequent differentiable normal map rendering.
We iteratively optimize the grid offsets $\mathcal{O}$ by minimizing the following objective function:
$\min_{\mathcal{O}}\sum\nolimits_{k=1}^{K}\left\|\mathcal{H}_{k}^{u}-\mathcal{H}_{k}\right\|_{2}^{2}+\omega_{s}\cdot\mathcal{L}_{s}+\omega_{n}\cdot\mathcal{L}_{n},$ (7)
where the first term aligns every pair of diffusion-predicted and mesh-rendered normal maps, the last two terms denote mesh vertex and normal regularizations weighted by factors $\omega_{s}$ and $\omega_{n}$. For each vertex $\mathbf{p}_{i}$ in the updated mesh and its neighboring vertices $\Omega_{i}=\{\mathbf{p}_{i,j}\}$, the Laplacian smoothing is computed by:
$\mathcal{L}_{s}=\sum\nolimits_{\mathbf{p}_{i}\in\mathcal{V}^{u}}\sum\nolimits_{\mathbf{p}_{i,j}\in\Omega_{i}}(\mathbf{p}_{i}-\mathbf{p}_{i,j})/|\Omega_{i}|,$ (8)
where $|\Omega_{i}|$ is the number of neighbors. Moreover, for each pair of neighboring normal vectors $\mathbf{n}_{i}$ and $\mathbf{n}_{i}^{\prime}$, we constrain local orientation consistency by:
$\mathcal{L}_{n}=\sum\nolimits_{i}1-(\mathbf{n}_{i}\cdot\mathbf{n}_{i}^{\prime})/(\left\|\mathbf{n}_{i}\right\|\cdot\left\|\mathbf{n}_{i}^{\prime}\right\|).$ (9)
After iterative optimization, we moderately apply Taubin smoothing [102] as post-processing to mitigate meaningless surface wrinkles.
4 Experiments
Figure 5: Visual comparison of normal map prediction results exported from (c) GeoWizard [27], (d) StableNormal [119], and (e) Ours, where (a) and (b) show the corresponding color images and ground-truth high-fidelity normal maps.
Table 1: Quantitative comparison of normal map prediction accuracy under the metric of mean angular error (MAE).
Method GeoWizard [27] StableNormal [119] Ours
MAE $\downarrow$ 17.59 14.33 9.05
Figure 6: The rendering effects of (c) our optimized mesh and (d) attaching the bump map obtained from PBRBoost [76], given the input textured coarse mesh as presented in (a) and (b).
Training Dataset Preparation. We used the publicly available Objaverse [23] dataset comprising 800K+ 3D assets to produce our training data. Through a series of meticulously designed filtering protocols and cleaning procedures, we ultimately acquired around 80K high-quality mesh models.
We used BlenderProc [24] for the rendering of color images, depth maps, and normal maps based on perspective projection. For each training mesh, we randomly sampled 16 viewpoints, with azimuth and elevation angles respectively varying within [0, 360] and [-30, 30].
To create paired low-poly and high-poly normal renderings, we performed geometry simplification on each high-poly training mesh model based on manifold reconstruction [36], and further removed detailed structures using [102]. We also manually annotated a smaller portion of unqualified training samples, which were used for training a separate neural model to further purify the whole training set.
Implementation Details. We adopted the popular diffusion architecture of Stable Diffusion V2.1 [87] and performed fine-tuning on 16 NVIDIA Tesla V100 GPUs for 40K steps with a total batch size of 128. We applied the AdamW [63] optimizer with cosine annealing, reducing the learning rate from 5e-5 to 1e-5. For both training and inference, the image resolution is configured to 512$\times$512.
For iterative mesh sculpting, we adopted Nvdiffrast [43] for efficient differentiable rendering from 12 camera poses (8 evenly-spaced horizontal views with azimuth angles from 0 to 360 and 4 downward-looking (elevated by 30) views with azimuth angles [45, 135, 225, 315]). Camera intrinsics were kept the same as our configurations in BlenderProc. The hyperparameters were empirically tuned as $r=512$, $\tau=0.5$, $\omega_{s}=0.25$, $\omega_{n}=0.01$ for 200 iterations.
4.1 Performance Evaluation
Figure 7: Comparison of mesh optimization effects, where (a) and (b) denote the input textured mesh, (c) and (d) present the results of our distance field deformation scheme and CR [76].
Table 2: Normal diffusion accuracy of different baseline schemes. Stoc. denotes the straightforward implementation of stochastic diffusion. w/o Sem. and w/o Geo. respectively denote removing the high-level semantics and low-level depth priors.
Method Original Stoc. w/o Sem. w/o Geo.
MAE $\downarrow$ 9.05 21.64 10.49 9.81
Figure 8: Illustration of prediction variation issues of the stochastic baseline diffusion scheme as displayed in (d). The color image, ground-truth normal map, and the output of our original design are presented in (a), (b), and (c).
Considering the lack of specialized research on our targeted task of texture-consistent 3D geometry super-resolution, we evaluated our superiority via applying SuperCarver on the textured meshes generated by various existing state-of-the-art 3D generation approaches, including InstantMesh [117], Unique3D [113], and CLAY [127], covering representative generation paradigms of feed-forward large reconstruction models, multi-view diffusion models, and native 3D latent diffusion models. As demonstrated in Figure 4, various input meshes are effectively boosted by our approach to produce realistic and expressive 3D surface details well aligned with the corresponding texture appearances.
Comparison on Normal Prediction. To evaluate the effectiveness of our specialized normal diffusion architecture, we made comparisons with advanced normal prediction frameworks, including GeoWizard [27] and StableNormal [119]. To facilitate quantitative evaluations, we manually selected 50 detail-rich textured meshes as ground-truths from Objaverse [23] and ABO [21] datasets. Then we applied surface simplification and smoothing to create geometrically-coarse meshes as the inputs.
As shown in Figure 5, our normal map predictions show better details and avoid obviously erroneous areas. Besides, we further deduced the metric of mean angular error (MAE) between ground-truth and predicted normal maps, where we masked invalid backgrounds for the outputs of GeoWizard and StableNormal. As illustrated in Table 1, owing to our specialized diffusion architecture, our normal map predictions significantly outperform the competing approaches. In addition, we also considered PBRBoost [107] (a latest material generation framework), since one of its outputs is the 2D bump maps associated with the base meshes, enabling us to compare its rendering effects when attaching the bump maps as height information. As shown in Figure 6, the actual fidelity of the enhanced geometric cues from PBRBoost is still far from satisfactory.
Comparison on Mesh Optimization. To facilitate surface geometry updating, many recent 3D generation frameworks [64, 20, 50, 113] introduce the CR [76] algorithm with joint optimization and remeshing, which is regarded as the most relevant competitor. In fact, we had previously attempted to adapt FlexiCubes [90], an advanced iso-surface representation, to achieve gradient-based geometry refinement starting from the initially given mesh surface. However, through extensive attempts, we empirically found that it easily leads to severe surface damages. We reason that this is attributable to its significantly increased deformation flexibility (which should have been its major strength) amplifies the impact of erroneous normal pixels. As shown in Figure 7, for typical hard cases with less accurate normal predictions, CR tends to produce noisy and degraded surfaces.
4.2 Ablation Study
To evaluate the effectiveness and necessity of the core components in our entire workflow, we conducted ablation study with different baseline schemes. We began with a straightforward baseline normal diffusion framework as formulated in Eq. (1) and Eq. (2). Besides, since prior guidance is also a critical condition for generating robust and accurate normal maps, we can design another two variants via removing the semantic cues extracted from DINO features or the geometric cues extracted from depth maps. As reported in Table 2, directly fine-tuning a standard diffusion model conditioned on color images lead to significantly degraded performance, with the MAE metric of 21.64. More importantly, as illustrated in Figure 8, the stochastic nature inevitably results in unstable predictions. As reported in the last two columns of Table 2, the joint injection of semantic and geometric priors are indeed beneficial. In the inverse normal rendering stage, we experimented with another two distance field deformation schemes, i.e., fixing grids while optimizing distances; jointly updating grids and distances. To obtain quantitative results, we computed the MAE metric of normal maps rendered from optimized meshes (note that Table 1 and Table 2 are targeted at diffused normal maps). As evidenced in Table 3, the two variants show sub-optimal performances, in that only optimizing distances usually causes weakened details while the joint optimization strategy easily suffers from surface distortion especially when the diffused normal maps contain relatively larger errors.
Table 3: Quantitative comparison of different distance field deformation schemes. Only Grids means our original design, Only Dist. and Joint G&D are the corresponding two baseline variants.
Method Only Grids Only Dist. Joint G&D
MAE $\downarrow$ 4.22 4.75 5.47
5 Conclusion
This paper presents a specialized geometry super-resolution framework for generating high-fidelity 3D surface details as depicted in the texture appearances of given coarse meshes, serving as a powerful tool for automated surface sculpting. Our framework is normal-centric and consists of two stages. The first stage predicts view-specific detail-boosted normal maps through a deterministic prior-guided diffusion architecture. The second stage transfers the underlying geometric patterns on multi-view normal map predictions onto the original mesh surface through a noise-resistant inverse rendering scheme driven by a carefully designed distance field deformation mechanism. Experiments demonstrate the superior performance of our SuperCarver across various shape types and texture styles. We believe that our exploration can make a difference to significantly accelerate the creation of high-poly meshes, and inspire more focused research in 3D geometry super-resolution, which is highly valuable yet basically ignored in the current community.
Limitations & Future Works. Our current framework still struggles to handle elongated objects (e.g., knives, swords) because the rendered areas are too small to provide adequate appearance information under finite image resolutions. One straightforward solution is to apply localized sculpting procedures. Besides, the inherent properties of signed distance field representations restrict our capabilities to handle open surfaces and complicated topologies. Hence, we might need to introduce differentiable iso-surface extractors built upon more generic implicit geometry representations.